If there are only Computer Science and Math Majors in Discrete Math, and there





are 4 Computer Science Majors, 15 Math Majors, and 8 student who are double





majors in Math and Computer Science, how many people are enrolled in discrete





math.

1
Expert's answer
2022-05-18T16:30:13-0400
"N(DM)=N(C\cup M)"

"=4+15+8=26"